Let y be 10% of a number. What is 50% of the number? Write your answer in terms of y.
Can you please show with process so I can understand it.

Found 2 solutions by jim_thompson5910, Theo:
Answer by jim_thompson5910(35256)

y = 10% of the number

y = 10% of x

y = 0.10x

We want 0.50x on the right side (to represent 50% of x), so multiply both sides by 5

5y = 5*0.10x

5y = 0.50x


The final answer is 5y

Answer by Theo(13342)

if y is equal to 10% of the number, then y * 5 = 50% of the number because 5 * 10% = 50%.

that's the logical interpretation.

the algebraic interpretation can be as follows:

y = .10 * x
solve for x to get x = y/.10

z = .50*x
replace x with y/.10 to get:
z = .50*y/.10
simplify to get:
z = 5*y

in the first equation, you're letting y = .10 * x
in the second equation, you're letting z = .50 * x
you solve for x in the first equation to get x = y/.10
you replace x with y/.10 in the second equation to get z = 5*y

the logical interpretation and the algebraic interpretation agree.
this is good.